# Matryoshka Representations for Adaptive Deployment

### v2

@@ Appendix E Image Retrieval
We carried out all retrieval experiments at $D_{s}$ $\in\{8,16,32,64,128,256,512,1024,2048\}$, as these were the representation sizes which were a part of the nesting_list at which losses were added during training, as seen in Algorithm 1, Appendix A. To examine whether $\rm MRL$} is able to learn $\rm Matryoshka~Representations$} at dimensions in between the representation size for which it was trained, we also tabulate the performance of $\rm MRL$} at interpolated $D_{s}$ $\in\{12,24,48,96,192,384,768,1536\}$ as $\rm MRL$}–Interpolated. We observed that performance scaled nearly monotonically between the original representation size and the interpolated representation size as we increase $D_{s}$, which demonstrates that $\rm MRL$} is able to learn $\rm Matryoshka~Representations$} at nearly all representation size $m\in[8,2048]$ despite optimizing only for $|\mathcal{M}|$ nested representation sizes.
@@ Appendix F Adaptive Retrieval
We also designed a simple cascade policy which we call funnel retrieval to successively improve and refine the k-NN shortlist at increasing $D_{s}$. This was an attempt to remove the dependence on manual choice of $D_{s}$ $D_{r}$. We retrieved a shortlist at $D_{s}$and then re-ranked the shortlist five times while simultaneously increasing $D_{r}$ (rerank cascade) and decreasing the shortlist length (shortlist cascade), which resembles a funnel structure. We tabulate the performance of funnel retrieval in various configurations in Table 13 on ImageNet-1K, and in Table 14 on ImageNet-4K. With funnel retrieval on ImageNet-1K, we were able to achieve top-1 accuracy within $0.1\%$ of retrieval with $D_{s}$ =2048=2048 (as in Table 8) with a funnel with $D_{s}$ =16=16, with $128\times$ less MFLOPs. Similarly, we are able to achieve equivalent top-1 accuracy within $0.15\%$ of retrieval at $D_{s}$ =2048=2048 (as in Table 10) with funnel retrieval at $D_{s}$ =32=32 on ImageNet-4K, with $64\times$ less MFLOPs. This demonstrates that with funnel retrieval, we can emulate the performance of retrieval with $D_{s}$ =2048=2048 with a fraction of the MFLOPs.
@@ Appendix G Few-shot and Sample Efficiency
We compare $\rm MRL$}, $\rm MRL\text{--}E$}, and FF  on various benchmarks to observe the effect of representation size on sample efficiency. We use Nearest Class Means [70] for classification which has been shown to be effective in the few-shot regime [13].
@@
We observe that $\rm MRL$} has equal performance to FF across all representation sizes and shot numbers. We also find that for both $\rm MRL$} and FF as the shot number decreases, the required representation size to reach optimal accuracy decreases (Table 15). For example, we observe that 1-shot performance at $32$ representation size has equal accuracy to $2048$ representation size.
Table 15: Few-shot accuracy (%) on ImageNetV2 for 1000 way classification. $\rm MRL$} performs equally to FF across all shots and representation sizes. We also observe that accuracy saturates at a lower dimension for lower shot numbers. For example, for 1-shot, 32-dim performs comparably to 2048-dim.
@@
For the long-tailed setting we evaluate $\rm MRL$} on the FLUID benchmark [83] which contains a mixture of pretrain and new classes. Table 16 shows the evaluation of the learned representation on FLUID. We observe that $\rm MRL$} provides up to 2% higher accuracy on novel classes in the tail of the distribution, without sacrificing accuracy on other classes. Additionally we find the accuracy between low-dimensional and high-dimensional representations is marginal for pretrain classes. For example, the 64-dimensional $\rm MRL$} performs $\sim 1\%$ lower in accuracy compared to the 2048-dimensional counterpart on pretrain-head classes (84.46% vs 85.60%). However for novel-tail classes the gap is far larger (6.22% vs 12.88%). We hypothesize that the higher-dimensional representations are required to differentiate the classes when few training examples of each are known. This results provides further evidence that different tasks require varying capacity based on their difficulty.
Table 16: Accuracy (%) categories indicates whether classes were present during ImageNet pretraining and head/tail indicates classes that have greater/less than 50 examples in the streaming test set. We observe that $\rm MRL$} performs better than the baseline on novel tail classes by $\sim 2\%$ on average.
@@ Appendix I In Practice Costs
Exact search has a search time complexity of $O(dkN)$, and HNSW has a search time complexity of $O(dk\log(N))$, where $N$ is the database size, $d$ is the representation size, and $k$ is the shortlist length. To examine real-world performance, we tabulate wall clock search time for every query in the ImageNet-1K and ImageNet-4K validation sets over all representation sizes $d$ in Table 19 for both Exact Search and HNSW32, and ablate wall clock query time over shortlist length $k$ on the ImageNet-1K validation set in Table 21. The wall clock time to build the index and the index size is also shown in Table 20.
@@
We analyzed the nature of model disagreement across representation sizes with $\rm MRL$} models with the help of Grad-CAM visualization [71]. We observed there were certain classes in ImageNet-1K such as "tools", "vegetables" and "meat cutting knife" which are occasionally located around multiple objects and a cluttered environment. In such scenarios, we observed that smaller representation size models would often get confused due to other objects and fail to extract the object of interest which generated the correct label. We also observed a different nature of disagreement arising when the models got confused within the same superclass. For example, ImageNet-1K has multiple "snake" classes, and models often confuse a snake image for an incorrect species of snake.
@@
We created a 30 superclass subset of the validation set based on wordnet hierarchy (Table 24) to quantify the performance of $\rm MRL$} model on ImageNet-1K superclasses. These 30 superclasses contain 467 out of 1000 classes, with an additional class “reject", when an image does not belong to any sueprclass. Table 25 quantifies the performance with different representation size. We observed that there was a jump in performance from 8 to 16 sized representations because predictions using first 8 dimensions was confusing images from the superclass "vegetable" with the reject token. We show 8 of these 30 superclasses and plot their top-1 accuracy (%) improvement with rep. size in Figure 11.
@@
Figure 11: Improvement of within-superclass classification with increasing representation size for several selected supervlasses. We observe that superclasses such as “oscine (songbird)” have a clear distinction between the object and background and thus predictions using representation size of 8 also lead to a good performance.
@@
Table 26: Top-1 classification accuracy (%) on ImageNet-1K of various ResNet50 models which are finetuned on pretrained FF-2048 model. We observed that adding more and more non-linearities is able to induce nesting to a reasonable extent even if the model was not pretrained with nesting in mind.
Rep. Size
@@
We performed an ablation of $\rm MRL$} over the relative importance, $c_{m}$, of different nesting dimensions $m\in\cal{M}$, as defined in Sec. 3. In an attempt to improve performance at lower dimensionalities, we boosted the relative importance $c_{m}$ of training loss at lower dimensions as in Eq. 1 with two models, $\rm MRL$}-8boost and $\rm MRL$}-8+16boost. The $\rm MRL$}-8boost model had $c_{m\in\cal M}=[2,1,1,1,1,1,1,1,1]$ and the $\rm MRL$}-8+16boost model had $c_{m\in\cal M}=[2,1.5,1,1,1,1,1,1,1]$. The relative importance list $c_{m\in\cal M}$ had a 1-to-1 correspondence with nesting dimension set $\mathcal{M}$. In Table 27, we observed that $\rm MRL$}-8boost improves top-1 accuracy by $3\%$ at $d=8$, and also improves top-1 of all representation scales from 16 to 256 over $\rm MRL$}, while only hurting the performance at 512 to 2048 representation scales by a maximum of 0.1%. This suggests that the relative importance $c_{m}$ can be tuned/set for optimal accuracy for all $m\in\mathcal{M}$, but we leave this extension for future work.
@@
To examine the effect of increasing shortlist lengths on search time, a reranking ablation over shortlist lengths is also performed for $D_{s}$= 16 and $D_{r}$= 2048 over ImageNet-1K  in Table 28, and over ImageNet-4K in Table 29. We observed that using a larger shortlist $k$ saturated ImageNet-1K performance at kk=200. But using larger shortlists until $k=2048$, the maximum value supported by the FAISS framework, steadily improved performance on ImageNet-4K. This is likely due to the increased database size, but could also indicate a correlation with ImageNet-4K being slightly out-of-distribution making the task at hand harder.

### v3

@@ Appendix E Image Retrieval
We carried out all retrieval experiments at $D_{s}$ $\in\{8,16,32,64,128,256,512,1024,2048\}$, as these were the representation sizes which were a part of the nesting_list at which losses were added during training, as seen in Algorithm 1, Appendix A. To examine whether $\rm MRL$} is able to learn $\rm Matryoshka~Representations$} at dimensions in between the representation size for which it was trained, we also tabulate the performance of $\rm MRL$} at interpolated $D_{s}$ $\in\{12,24,48,96,192,384,768,1536\}$ as $\rm MRL$}–Interpolated and $\rm MRL\text{--}E$}–Interpolated (see Table 8). We observed that performance scaled nearly monotonically between the original representation size and the interpolated representation size as we increase $D_{s}$, which demonstrates that $\rm MRL$} is able to learn $\rm Matryoshka~Representations$} at nearly all representation size $m\in[8,2048]$ despite optimizing only for $|\mathcal{M}|$ nested representation sizes.
@@ Appendix F Adaptive Retrieval
We also designed a simple cascade policy which we call funnel retrieval to successively improve and refine the k-NN shortlist at increasing $D_{s}$. This was an attempt to remove the dependence on manual choice of $D_{s}$ & $D_{r}$. We retrieved a shortlist at $D_{s}$ and then re-ranked the shortlist five times while simultaneously increasing $D_{r}$ (rerank cascade) and decreasing the shortlist length (shortlist cascade), which resembles a funnel structure. We tabulate the performance of funnel retrieval in various configurations in Table 13 on ImageNet-1K, and in Table 14 on ImageNet-4K. With funnel retrieval on ImageNet-1K, we were able to achieve top-1 accuracy within $0.1\%$ of retrieval with $D_{s}$ =2048=2048 (as in Table 8) with a funnel with $D_{s}$ =16=16, with $128\times$ less MFLOPs. Similarly, we are able to achieve equivalent top-1 accuracy within $0.15\%$ of retrieval at $D_{s}$ =2048=2048 (as in Table 10) with funnel retrieval at $D_{s}$ =32=32 on ImageNet-4K, with $64\times$ less MFLOPs. This demonstrates that with funnel retrieval, we can emulate the performance of retrieval with $D_{s}$ =2048=2048 with a fraction of the MFLOPs.
@@ Appendix G Few-shot and Sample Efficiency
We compared $\rm MRL$}, $\rm MRL\text{--}E$}, and FF on various benchmarks to observe the effect of representation size on sample efficiency. We used Nearest Class Means [74] for classification which has been shown to be effective in the few-shot regime [13].
@@
We observed that $\rm MRL$} had equal performance to FF across all representation sizes and shot numbers. We also found that for both $\rm MRL$} and FF, as the shot number decreased, the required representation size to reach optimal accuracy decreased (Table 15). For example, we observed that 1-shot performance at $32$ representation size had equal accuracy to $2048$ representation size.
Table 15: Few-shot accuracy (%) on ImageNetV2 for 1000-way classification. $\rm MRL$} performs equally to FF across all shots and representation sizes. We also observed that accuracy saturated at a lower dimension for lower shot numbers. E.g. for 1-shot, 32-dim performed comparably to 2048-dim.
@@
For the long-tailed setting we evaluated $\rm MRL$} on the FLUID benchmark [87] which contains a mixture of pretrain and new classes. Table 16 shows the evaluation of the learned representation on FLUID. We observed that $\rm MRL$} provided up to 2% higher accuracy on novel classes in the tail of the distribution, without sacrificing accuracy on other classes. Additionally we found the accuracy between low-dimensional and high-dimensional representations was marginal for pretrain classes. For example, the 64-dimensional $\rm MRL$} performed $\sim 1\%$ lower in accuracy compared to the 2048-dimensional counterpart on pretrain-head classes (84.46% vs 85.60%). However for novel-tail classes the gap was far larger (6.22% vs 12.88%). We hypothesize that the higher-dimensional representations are required to differentiate the classes when few training examples of each are known. These results provide further evidence that different tasks require varying capacity based on their difficulty.
Table 16: Accuracy (%) categories indicates whether classes were present during ImageNet pretraining and head/tail indicates classes that have greater/less than 50 examples in the streaming test set. We observed that $\rm MRL$} performed better than the baseline on novel tail classes by $\sim 2\%$ on average.
@@ Appendix I In Practice Costs
Exact search has a search time complexity of $O(dkN)$, and HNSW has a search time complexity of $O(dk\log(N))$, where $N$ is the database size, $d$ is the representation size, and $k$ is the shortlist length. To examine real-world performance, we tabulated wall clock search time for every query in the ImageNet-1K and ImageNet-4K validation sets over all representation sizes $d$ in Table 19 for both Exact Search and HNSW32, and ablated wall clock query time over shortlist length $k$ on the ImageNet-1K validation set in Table 21. The wall clock time to build the index and the index size is also shown in Table 20.
@@
We analyzed the nature of model disagreement across representation sizes with $\rm MRL$} models with the help of Grad-CAM visualization [75]. We observed there were certain classes in ImageNet-1K such as "tools", "vegetables" and "meat cutting knife" which were occasionally located around multiple objects and a cluttered environment. In such scenarios, we observed that smaller representation size models would often get confused due to other objects and fail to extract the object of interest which generated the correct label. We also observed a different nature of disagreement arising when the models got confused within the same superclass. For example, ImageNet-1K has multiple "snake" classes, and models often confuse a snake image for an incorrect species of snake.
@@
We created a 30 superclass subset of the validation set based on wordnet hierarchy (Table 24) to quantify the performance of $\rm MRL$} model on ImageNet-1K superclasses. Table 25 quantifies the performance with different representation size.
@@
Table 26: Top-1 classification accuracy (%) on ImageNet-1K of various ResNet50 models which are finetuned on pretrained FF-2048 model. We observed that adding more non-linearities is able to induce nesting to a reasonable extent even if the model was not pretrained with nesting in mind.
Rep. Size fc
@@
We performed an ablation of $\rm MRL$} over the relative importance, $c_{m}$, of different nesting dimensions $m\in\cal{M}$, as defined in Sec. 3. In an attempt to improve performance at lower dimensionalities, we boosted the relative importance $c_{m}$ of training loss at lower dimensions as in Eq. 1 with two models, $\rm MRL$}-8boost and $\rm MRL$}-8+16boost. The $\rm MRL$}-8boost model had $c_{m\in\cal M}=[2,1,1,1,1,1,1,1,1]$ and the $\rm MRL$}-8+16boost model had $c_{m\in\cal M}=[2,1.5,1,1,1,1,1,1,1]$. The relative importance list $c_{m\in\cal M}$ had a 1-to-1 correspondence with nesting dimension set $\mathcal{M}$. In Table 27, we observed that $\rm MRL$}-8boost improves top-1 accuracy by $3\%$ at $d=8$, and also improves top-1 accuracy of all representation scales from 16 to 256 over $\rm MRL$}, while only hurting the performance at 512 to 2048 representation scales by a maximum of 0.1%. This suggests that the relative importance $c_{m}$ can be tuned/set for optimal accuracy for all $m\in\mathcal{M}$, but we leave this extension for future work.
$\rm Matryoshka~Representations$} at Arbitrary Granularities.
To train $\rm MRL$}, we used nested dimensions at logarithmic granularities $\mathcal{M}=\{8,16,\ldots,1024,2048\}$ as detailed in Section 3. We made this choice for two empirically-driven reasons: a) The accuracy improvement with increasing representation size was more logarithmic than linear (as shown by FF models in Figure 3). This indicated that optimizing for granularities increasing in a non-logarithmic fashion would be sub-optimal both for maximum performance and expected efficiency; b) If we have $m$ arbitrary granularities, the expected cost of the linear classifier to train $\rm MRL$} scales as $O(L*(m^{2}))$ while logarithmic granularities result in O(L2log(d))O(L*2log(d)) space and compute costs.
To demonstrate this effect, we learned $\rm Matryoshka~Representations$} with uniform ($\rm MRL$}-Uniform) nesting dimensions $m\in\mathcal{M}=\{8,212,416,620,824,1028,1232,1436,1640,1844,2048\}$. We evaluated this model at the standard ($\rm MRL$}-log) dimensions $m\in\mathcal{M}=\{8,16,32,64,128,256,512,1024,2048\}$ for ease of comparison to reported numbers using 1-NN accuracy (%). As shown in Table 29, we observed that while performance interpolated, $\rm MRL$}-Uniform suffered at low dimensions as the logarithmic spacing of $\rm MRL$}-log resulted in tighter packing of information in these initial dimensions. The higher nesting dimensions of $\rm MRL$}-Uniform did not help in significant accuracy improvement due to accuracy saturation, which is often logarithmic in representation size as shown by FF models. Note that the slight improvement at dimensions higher than 512 for $\rm MRL$}-Uniform is due to multiple granularities around them compared to just three for $\rm MRL$}-log, which are not useful in practice for efficiency.
Lower Dimensionality.
We experimented with training $\rm MRL$} with smaller nesting dimension than $m=8$, as shown in Table 29, with two models: MRL-4 and MRL-6. We found that using lower than 8-dimensions to train $\rm MRL$}, i.e. $m_{0}\in\{4,6\}$ for MRL-4 and MRL-6 respectively, did not affect the top-1 accuracy of other granularities significantly. However, granularities smaller than 8-dimensions had very low accuracy and were often unusable for deployment along with additional training difficulty. We also observed a small dip in accuracy at higher dimensions which we attribute to the joint loss that now also included the harder optimization of the smallest dimension. Lastly, we hypothesize the dimensionality of 8 is an empirically validated design choice due to the considerable accuracy it provided along with the ease of training.
Table 28: An ablation over training with smaller nesting dimensionalities in terms of Top-1 accuracy (%). MRL-4 and MRL-6 are variations of the original model (MRL-8) with $m_{0}\in\{4,6\}$, where $m\in\mathcal{M}$ is part of the nesting_list as seen in Alg 2.
Rep. Size MRL-4 MRL-6 MRL-8
4 27.25 - -
6 - 58.71 -
8 66.86 67.55 66.63
16 73.36 73.10 73.53
32 74.82 74.49 75.03
64 75.51 75.32 75.82
128 75.93 75.61 76.30
256 76.08 75.82 76.47
512 76.31 75.93 76.65
1024 76.38 76.04 76.76
2048 76.43 76.12 76.80
Table 29: An ablation over training $\rm MRL$} with nesting list at uniformly distributed granularities. Entries in the $\rm MRL$}-Uniform column are evaluated at logarithmic dimensions for a fair comparison to $\rm MRL$}-Log (standard $\rm MRL$}) with 1-NN accuracy (%).
Rep. Size $\rm MRL$}-Log $\rm MRL$}-Uniform
8 62.19 58.44
16 67.91 61.11
32 69.46 63.82
64 70.17 66.44
128 70.52 68.71
256 70.62 70.06
512 70.82 70.98
1024 70.89 71.37
2048 70.97 71.44
@@
To examine the effect of increasing shortlist lengths on search time, we performed a reranking ablation over shortlist lengths for $D_{s}$= 16 and $D_{r}$= 2048 over ImageNet-1K  in Table 30, and over ImageNet-4K in Table 31. We observed that using a larger shortlist $k$ saturated ImageNet-1K performance at kk=200. But using larger shortlists until $k=2048$, the maximum value supported by the FAISS framework, steadily improved performance on ImageNet-4K. This is likely due to the increased database size, but could also indicate a correlation with ImageNet-4K being slightly out-of-distribution making the task at hand harder.